The 2 roots of ax^2+bx+c=0 are both 2 digit integers with the same units digit, but D/a^2 (D=Discriminant) is not a multiple of 100. Which is a possible value of the unit digit of b/2a?

a)4, b)5, c)6, d)7

Grade:Upto college level

2 Answers

Ramesh V
14 years ago

Let the 2 roots be (10a+k) and (10b+k)

here k is unit digit of both roots

so sum of roots = -b/a

product of roots = c/a

(10(a+b) +2k) = -b/a    and     (10a+k)*(10b+k) = c/a

k = -b/2a -5(a+b) which on substitution gives :

(5(a-b) - b/2a)*(5(b-a) + b/2a) = c/a

(b/2a - 5(a-b))*(5(a-b) + b/2a) = c/a

b2/4a2 - 25(a-b)2 = c/a

(b2 - 4ac)/a2 = 100(a-b)2

but given that (b2 - 4ac)/a2 is not multiple of 100 [ as a,b are intezers, a-b

so (a-b)2<1

i.e., possible value of :  b/2a =1/2

 

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch

dhirendra pratap singh
14 years ago

let the roots are 10p and 10q.

now,

sum of roots: 10(p+q)= -b/a.

product of roots is 100pq= c/a

value of  d/a^2 = 100(p-q)^2.

since d/a^2 is not a multiple of 100, hence (p-q) is not integer.

thus p and q are not integers.

but for roots being integers p and q should be multiple of 1/ 2  or 1/ 5.

let both are multiple of 1/2

then both roots must have 0 or 5 at its unit place.

thus the possible value of unit digit of b/2a is 0 or 5.

hence  the correct option is b.